Appling Enterprises issued 9% bonds with a face amount of $510,000 on January 1, 2018. The bonds sold for $466,244 and mature in 2037 (20 years). For bonds of similar risk and maturity the market yield was 10%. Interest is paid semiannually on June 30 and December 31. Appling determines interest expense at the effective rate. Appling elected the option to report these bonds at their fair value. The fair values of the bonds at the end of each quarter during 2018 as determined by their market values in the over-the-counter market were the following:

March 31 $490,000
June 30 470,000
September 30 465,000
December 31 472,000

   
General (risk-free) interest rates did not change during 2018.

Required:

1. By how much will Appling’s comprehensive income be increased or decreased by the bonds (ignoring taxes) in the March 31 quarterly financial statements?
2. By how much will Appling’s comprehensive income be increased or decreased by the bonds (ignoring taxes) in the June 30 quarterly financial statements?
3. By how much will Appling’s comprehensive income be increased or decreased by the bonds (ignoring taxes) in the September 30 quarterly financial statements?
4.By how much will Appling’s comprehensive income be increased or decreased by the bonds (ignoring taxes) in the December 31 annual financial statements?
(For all requirements, Do not round your intermediate calculations.)
Answer #1

Part 1

March 31 book value (amortized initial amount) = 466244 + (5%*466244*3/6) = 477900

Fair value adjustment = 477900-490000 = (12100)

Decrease in earnings = 11656+121000 = 132656

Part 2

June 30 book value (amortized initial amount) = 477900-(510000*4.5%*3/6)+((5%*466244*3/6)-(510000*4.5%*3/6)) = 466606

Fair value adjustment = 466606-470,000 =-3394

Change in fair value adjustment =-3394-(-12100) = 8706

Decrease in earnings = (5%*466244*3/6)-8706 = 2950

Part 3

September 30 book value (amortized initial amount) = 466606+(510000*4.5%*3/6)+((5%*((466244+181+181)*3/6)-(510000*4.5%*3/6)) = 478271

Fair value adjustment balance =478271-465,000 = 13271

Change in fair value adjustment = 13271-(-3394) = 16665

Increase in earnings = 16665-((5%*((466244+181+181)*3/6)) = 5000

Part 4

December 31 book value = 478271-(510000*4.5%*3/6)+((5%*((466244+181+181)*3/6)-(510000*4.5%*3/6)) = 466986

Fair value adjustment = 466986-472,000 = -5014

Change in fair value adjustment = -5014-13271 = -18285

Decrease in earnings = 11656+11656+11665+11665+12100-8706-16665+18285 = 51656